13.2g of a gas occupies a volume of 0.918dm³ at 25°C and a pressure of 900 torr was found to have 2.73 g of a gas calculate the molecular mass of the gas?

1
Expert's answer
2021-11-17T13:28:03-0500

"m=13.2 \;g \\

V = 0.918 \; dm^3 = 0.918 \; L \\

T = 25 + 273.15 = 298.15 \;K \\

p = 900 \; torr = 1.184 \;atm"

Ideal gas law

pV =nRT

"R= 0.08206 \; L×atm/mol×K \\

n = \frac{pV}{RT} \\

n = \frac{1.184 \times 0.918}{0.08206 \times 298.15} = 0.0444 \; mol \\

M = \frac{m}{n} \\

M = \frac{13.2}{0.0444} = 297.29 \; g/mol"

Answer: 297.29 g/mol